# Residually finite groups with uniformly almost flat quotients

### v1

@@ 8. Polycyclic groups
In the later subsections, we will make use the maximal nilpotent normal subgroup of a polycyclic group. Recall that the Fitting subgroup of a group $G$ is the subgroup generated by all the nilpotent normal subgroups, i.e. $\operatorname{Fit}(G)=\langle N|N\trianglelefteqslant G,\text{$N$ nilpotent}\rangle$. For a polycyclic group $P$, the Fitting subgroup $\operatorname{Fit}(P)$ is nilpotent [24, Corollary 13 on p. 15].
@@ 8.2. Some linear algebra results
Here are some useful equivalent characterization of matrices in $\operatorname{GL}(m,\mathbb{Z})$ from[21, Section 4].
@@
The following theorem is a special case of [24, Exercise 7 on p. 92].
@@ 8.4. Groups of the form $P=\mathbb{Z}^{n}\rtimes\mathbb{Z}^{m}$
The following result from [29, Proposition 2.2] gives an nice sufficient condition for two semidirect products to be isomorphic.

### v2

@@ 8. Polycyclic groups
In the later subsections, we will make use the maximal nilpotent normal subgroup of a polycyclic group. Recall that the Fitting subgroup of a group $G$ is the subgroup generated by all the nilpotent normal subgroups, i.e. $\operatorname{Fit}(G)=\langle N|N\trianglelefteqslant G,\text{$N$ nilpotent}\rangle$. For a polycyclic group $P$, the Fitting subgroup $\operatorname{Fit}(P)$ is nilpotent [30, Corollary 13 on p. 15].
@@ 8.2. Some linear algebra results
Here are some useful equivalent characterization of matrices in $\operatorname{GL}(m,\mathbb{Z})$ from[27, Section 4].
@@
The following theorem is a special case of [30, Exercise 7 on p. 92].
@@ 8.4. Groups of the form $P=\mathbb{Z}^{n}\rtimes\mathbb{Z}^{m}$
The following result from [36, Proposition 2.2] gives an nice sufficient condition for two semidirect products to be isomorphic.